# Simulation of Hypergraph Algorithms with Looped Transformers

### v1

@@ 4 Key Operation Implementation
In Section 4.1, the operation of selection is discussed. In Section 4.2, the operation of increment is discussed. In Section 4.3, the operation of reading scalar from column is discussed. In Section 4.4, the operation of comparison is discussed. In Section 4.5, the operation of writing scalar to column is discussed. In Section 4.6, the operation of termination is discussed. In Section 4.7, the operation of reading from incident matrix is discussed. In Section 4.8, the operation of AND is discussed. In Section 4.9, the operation of repeat AND is discussed. In Section 4.10, the operation of repeat addition is discussed.
4.1 Selection
Lemma 4.1 (Selection).
@@ 4.1 Selection
4.2 Increment
Lemma 4.2 (Increment).
@@ 4.2 Increment
  • Let $C_{1},C_{2}$ be the index for source position embedding, $D$ be the index for target position embedding.
@@
4.3 Read Scalar From Column
Lemma 4.3 (Read scalar from column).
@@ 4.3 Read Scalar From Column
  • Let $C$ be the index for source row, $D$ be the index for source column.
@@
Algorithm 3 Iteration of hyperedge
1:$\mathrm{idx}_{\mathrm{hyperedge}}\leftarrow 0$
[…]
3:$\mathrm{visit}_{\mathrm{hyperedge}},~{}\mathrm{candidates}_{\mathrm{hyperedge}}\leftarrow{\bf 0}_{n_{e}}\|{\bf 1}_{K-1-n_{e}},~{}\Omega\cdot{\bf 1}_{K-1}$
4:$\mathrm{termination}_{\mathrm{hyperedge}}\leftarrow 0$triangleright Initialization of visiting hyperedges
5:
6:procedure VisitHyperedge($\widetilde{A}\in\mathbb{R}_{+}^{n_{v}\times n_{e}}$)
7:     if $\mathrm{termination}_{\mathrm{hyperedge}}=1$ then
8:         $\cdots$ $\triangleright$ Re-initialization of visiting edge, Lemma 4.1
9:         $\mathrm{termination}_{\min}\leftarrow 0$
10:     end if
11:     $\mathrm{idx}_{\mathrm{hyperedge}}\leftarrow\mathrm{idx}_{\mathrm{hyperedge}}+1$triangleright Increment, Lemma 4.2
12:     $\mathrm{val}_{\mathrm{hyperedge}}\leftarrow A[:,\mathrm{idx}_{\mathrm{hyperedge}}]$triangleright Read column from incident matrix, Lemma 4.7
13:     for $i=1\to K-1$ do
14:         $\mathrm{iszero}_{\mathrm{hyperedge}}[i]\leftarrow(\mathrm{val}_{\mathrm{hyperedge}}[i]\leq 0)$triangleright Compare, Lemma 4.4
15:     end for
16:     for $i=1\to K-1$ do
17:         if $\mathrm{iszero}_{\mathrm{hyperedge}}[i]=1$  then
18:              $\mathrm{val}_{\mathrm{hyperedge}}[i]\leftarrow\Omega$triangleright Selection, Lemma 4.1
19:         end if
20:     end for
21:     for $i=1\to K-1$ do
22:         if $\mathrm{val}_{\mathrm{hyperedge}}[i]<\mathrm{candidates}[i]$ then $\triangleright$ Compare, Lemma 4.4
23:              $\mathrm{candidates}[i]\leftarrow\mathrm{val}_{\mathrm{hyperedge}}[i]$triangleright Update variables, Lemma 4.1
24:         end if
25:     end for
26:     $\mathrm{visit}_{\mathrm{hyperedge}}[\mathrm{idx}_{\mathrm{hyperedge}}]\leftarrow 1$triangleright Write scalar to column, Lemma 4.5
27:     $\mathrm{termination}_{\mathrm{hyperedge}}\leftarrow\neg(0~{}\mathrm{in}~{}\mathrm{visit}_{\min})$triangleright Trigger termination, Lemma 4.6
28:     $\mathrm{termination}_{\min}\leftarrow\mathrm{termination}_{\min}\land\mathrm{termination}_{\mathrm{hyperedge}}$triangleright AND, Lemma 4.8
29:end procedure
Algorithm 4 Dijkstra’s algorithm for shortest path
5:
6:    while $\mathrm{termination}$ is $\mathbf{false}$ do
7:         for $i=1\to K-1$ do
8:             if $\mathrm{visit}[i]$ is $\mathbf{true}$ then
9:                 $\mathrm{dists}_{\mathrm{masked}}[i]\leftarrow\Omega$triangleright Selection, Lemma 4.1
10:             else
11:                 $\mathrm{dists}_{\mathrm{masked}}[i]\leftarrow\mathrm{dists}[i]$
12:             end if
13:         end for
14:
15:         GetMinimumValue $(\mathrm{dists}_{\mathrm{masked}})$triangleright Get minimum value, Lemma 5.2
16:
17:         if $\mathrm{termination}_{\min}$ is $\mathbf{true}$ then
18:             $\mathrm{node}\leftarrow\mathrm{idx}_{\mathrm{best}}$
19:             $\operatorname{dist}\leftarrow\mathrm{val}_{\mathrm{best}}$triangleright Selection, Lemma 4.1
20:         end if
21:
22:         VisitHyperdege $(A)$triangleright Degradation, Theorem 6.1
23:
24:         if $\mathrm{termination}_{\mathrm{hyperedge}}=1$ then
25:             $\mathrm{candidates}\leftarrow\mathrm{candidates}_{\mathrm{hyperedge}}$triangleright Selection, Lemma 4.1
26:         end if
27:         for $i=1\to K-1$ do
28:             $\mathrm{candidates}[i]\leftarrow\mathrm{candidates}[i]+\operatorname{dist}$triangleright Addition, Lemma 4.10
29:         end for
30:         for $i=1\to K-1$ do
31:             $\mathrm{changes}[i]\leftarrow\mathrm{candidates}[i]<\mathrm{dists}[i]$triangleright Compare, Lemma 4.4
32:         end for
33:         for $i=1\to K-1$ do
34:             if $\mathrm{termination}_{\min}=0$ and $\mathrm{iszero}[i]=1$ then
35:                 $\mathrm{changes}[i]\leftarrow 0$triangleright repeat AND, Lemma 4.9
36:             end if
37:         end for
38:         for $i=1\to K-1$ do
39:             if $\mathrm{changes}[i]=1$ then
40:                 $\mathrm{prev}[i],~{}\mathrm{dists}[i]\leftarrow\mathrm{node},\mathrm{candidates}[i]$triangleright Selection. Lemma 4.1
41:             end if
42:         end for
43:         $\mathrm{visit}[\mathrm{node}]\leftarrow\mathrm{visit}[\mathrm{node}]+\mathrm{termination}_{\min}$triangleright Write scalar to column, Lemma 4.5
44:         $\mathrm{termination}\leftarrow\neg(0~{}\mathrm{in}~{}\mathrm{visit})$triangleright Trigger termination, Lemma 4.6
45:    end while
46:    return $\mathrm{prev},~{}\mathrm{dists}$
47:end procedure
Algorithm 5 Helly, Algorithm 3 in [5]
1:$\mathrm{idx}_{x}\leftarrow 0$
2:$\mathrm{idx}_{y}\leftarrow 1$
3:$\mathrm{idx}_{v}\leftarrow 0$
4:$\mathrm{termination}\leftarrow 0$
5:$\mathrm{helly}\leftarrow 1$triangleright Initialization of Helly
6:
7:while $\mathrm{termination}=0$ do
8:     $\mathrm{hyperedge}_{x}\leftarrow A[:,\mathrm{idx}_{x}]$
9:     $\mathrm{hyperedge}_{y}\leftarrow A[:,\mathrm{idx}_{y}]$
10:     $\mathrm{hyperedge}_{v}\leftarrow A[:,\mathrm{idx}_{v}]$triangleright Read column from incident matrix, Lemma 4.7
11:     $\mathrm{hyperedge}_{x}\leftarrow\mathrm{hyperedge}_{x}>0$
12:     $\mathrm{hyperedge}_{y}\leftarrow\mathrm{hyperedge}_{y}>0$
13:     $\mathrm{hyperedge}_{v}\leftarrow\mathrm{hyperedge}_{v}>0$triangleright Compare, Lemma 4.4
14:     $\mathrm{intersection}\leftarrow\mathrm{hyperedge}_{x}\wedge\mathrm{hyperedge}_{y}\wedge\mathrm{hyperedge}_{v}$triangleright AND, Lemma 4.8
15:     $\mathrm{helly}_{v}\leftarrow\neg$(1 in $\mathrm{intersection}$)
16:     if $\mathrm{helly}_{v}=1$ then
17:         $\mathrm{helly}\leftarrow 0$
18:         $\mathrm{termination}\leftarrow 1$triangleright Selection, Lemma 4.1
19:     end if
20:     $\mathrm{idx}_{v}\leftarrow\mathrm{idx}_{v}+1$triangleright Increment, Lemma 4.2
21:     if $\mathrm{idx}_{v}>n_{v}$ then
22:         $\mathrm{idx}_{v}\leftarrow 1$triangleright Selection, Lemma 4.1
23:         $\mathrm{idx}_{y}\leftarrow\mathrm{idx}_{y}+1$triangleright Selection and Increment, Lemma 4.1 and Lemma 4.2
24:     end if
25:     if $\mathrm{idx}_{y}>n_{v}$ then
26:         $\mathrm{idx}_{x}\leftarrow\mathrm{idx}_{x}+1$
27:         $\mathrm{idx}_{y}\leftarrow\mathrm{idx}_{x}+1$triangleright Selection and Increment, Lemma 4.1 and Lemma 4.2
28:     end if
29:     if $\mathrm{idx}_{x}=n_{v}$ then
30:         $\mathrm{termination}\leftarrow 1$triangleright Selection, Lemma 4.1
31:     end if
32:end while
33:return $\mathrm{helly}$
4.4 Comparison
Lemma 4.4 (Comparison).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.11.
  • Let $\Omega$ represent the maximum absolute value within a clause.
  • Let $C,~{}D$ be the index for the column for comparing.
  • Let $E$ be the index for the target column to write the comparison result.
Proof.
4.5 Write Scalar to Column
Lemma 4.5 (Write scalar to column).
@@ 4.5 Write Scalar to Column
4.6 Termination
Lemma 4.6 (Termination).
@@ 4.6 Termination
4.7 Read from Incident Matrix
Lemma 4.7 (Read from incident matrix).
@@ 4.8 AND
4.9 Repeat AND
Lemma 4.9 (Repeat AND).
@@ 4.9 Repeat AND
4.10 Repeat Addition
Lemma 4.10 (Repeat addition).
@@ 4.10 Repeat Addition
5 Simulation
We present the simulation result of visiting hyperedges iteratively in Section 5.1. We discuss the result of getting a minimum value in Section 5.2. We discuss the simulation result of Dijkstra’s Algorithm in Section 5.3.
5.1 Iteration of Visiting Hyperedges
Lemma 5.1 (Visiting hyperedges iteratively).

### v2

@@ 5 Key Operation Implementation
In this section, we introduce some basic operations that can be simulated by a single-layer transformer, which serve as fundamental primitives in our theoretical framework. Using these primitives, we can simulate complex deterministic algorithms that use hypergraphs or hyperedges as iterative objects.
Selection.
Here, we present the selection operation. Given a boolean array, referred to as condition $C$, a target array $E$, and two value arrays, $V_{0}$ and $V_{1}$, the selection operation can achieve the following: when the condition $C$ is $1$, the value from $V_{1}$ will be written to the target array $E$, when the condition $C$ is 00, the value from $V_{0}$ will be written to the target array $E$. By this operation, we can combine two arrays into a single array. Furthermore, by setting the target array $E$ to be the same as one of the value arrays, i.e., $V_{0}$ or $V_{1}$, we can realize conditional updates. This means that the value of an array will be updated if and only if a certain condition holds.
If $V_{0}$, $V_{1}$, $E$, and $C$ are scalars, this operation can be directly applied to scalars, as the lower $[K-1]$ values are kept as 00. In the following lemma, we aim to show that the selection operation can be simulated by a single-layer transformer. See details in Lemma C.1.
Increment.
Here, we present the operation of increment using positional embeddings. Let $C_{1}$ and $C_{2}$ represent the positional embeddings corresponding to $\sin(\cdot)$ and $\cos(\cdot)$, respectively. The primary objective is to employ a rotation matrix to update the angles encoded in $C_{1}$ and $C_{2}$ by increasing the angular offset $\widehat{\delta}$. The updated values are then written to the target locations $D_{1}$ and $D_{2}$. Typically, when $C_{1}=D_{1}$ and $C_{2}=D_{2}$, this operation can be performed in-place. See details in Lemma C.2.
Comparason.
Here, we introduce the operation of comparison, which involves comparing two arrays. In greedy algorithms, conditional updates are often employed, where the stored optimal solution is updated only if the current solution is better than the previously stored one. This process is captured in the condition update within the selection operation, as described in Lemma C.1. Notably, this requires a Boolean array as the condition input. For example, in Dijkstra’s algorithm, we evaluate whether a path is shorter and update only those paths that are shorter than the currently stored shortest paths. See details in Lemma C.3.
Read Scalar From Column.
We present the operation of reading a scalar from a column. Let the target column be denoted as $E$, the position embeddings of the source row $C$ as $C_{1}$ and $C_{2}$, and the source column as $D$. The array is stored in column $D$, and our objective is to extract the scalar at the $C$-th position and write it to the top row of column $E$. In an MLP layer represented as $XW$, we can extract a column using its column index. However, to extract a specific row index, we leverage the property of the positional embedding, $p_{i}^{\top}p_{j}<p_{i}^{\top}p_{i}$ for $i\neq j$. This operation enables more flexible operations on individual values within the array. See details in Lemma C.4.
Write Scalar to Column.
We present the operation of writing a scalar to a column. This operation parallels the process of reading a scalar from a column. The construction in this step also employs position embedding, analogous to Lemma C.4. See details in Lemma C.5.
Termination.
In the greedy algorithm, we terminate the process and return the result after traversing all objects. Here, we maintain an array, marking the traversed objects as 1. Notably, since the value of $K-1$ does not always equal the number of objects (typically $n_{v}$ or $n_{e}$), we fill the remaining positions with 1. See details in Lemma C.6.
Read from Incident Matrix.
Instead of storing a static incidence matrix in a matrix $X\in\mathbb{R}^{K\times d}$, we utilize an attention-head-incorporated incidence matrix as defined in Definition 3.6. This construction allows $d$ to be a constant independent of $n_{e}$ and $n_{v}$, implying that the feature dimension of the model $h_{T}$ can be controlled in $O(1)$. See details in Lemma C.7.
AND.
This operation can be applied to both arrays and scalars because, even when the input column is always 00, the result of the AND operation will remain 00. This operation is particularly useful when combining two conditions to trigger a specific operation. See details in Lemma C.8.
Repeat AND.
Different from the regular AND operation described in Lemma C.8, the repeat AND operation combines a scalar and an array. It can be understood as first replicating the scalar into a length-$K-1$ array, followed by performing the regular AND operation. This operation is often used in nested if statements. See details in Lemma C.9.
Repeat Addition.
Similar to the repeat AND operation, the repeat addition operation first replicates a scalar into an array and then performs element-wise addition between two arrays. This operation is commonly used when updating the distance to the next vertices. See details in Lemma C.10.
6 Simulation
We present the simulation result of visiting hyperedges iteratively in Section 6.1. We discuss the simulation result of Dijkstra’s Algorithm in Section 6.2.
6.1 Iteration of Visiting Hyperedges
First, we present our result on visiting hyperedges iteratively. See details in Algorithm 2.
Lemma 6.1 (Visiting hyperedges iteratively, informal version of Lemma D.2).
6.2 Dijkstra’s Algorithm
Furthermore, we combine the iteratively visiting hyperedge pattern with Dijkstra’s Algorithm to extend it to a hypergraph. For details, see Algorithm 4.
Theorem 6.2 (Dijkstra’s Algorithm on hypergraph, informal version of Theorem D.1).
7 Conclusion
Appendix
Roadmap. In Section A, we introduce more related work. In Section B, we introduce some tools from previous work. In Section C, we present the missing lemma in Section 5. In Section D, we present the missing proof in Section 6. In Section E, we present the missing proof in Section 4.
Appendix A More Related Work
A.1 Large Language Models
Transformer-based neural networks [97] have rapidly become the leading framework for natural language processing within machine learning. When these models scale to billions of parameters and are trained on expansive, heterogeneous data, they are frequently called large language models (LLMs) or foundation models [8]. Representative LLMs include BERT [21], PaLM [19], Llama [93], ChatGPT [76], and GPT4 [74]. Such models exhibit versatile capabilities [5] across a broad array of downstream tasks.
In pursuit of optimizing LLMs for specific applications, a variety of adaptation strategies have been introduced. These range from using adapters [45, 113, 32, 82], calibration methods [118, 115], and multitask fine-tuning [30, 106, 96, 107], to prompt tuning [31, 53], scratchpad techniques [73], instruction tuning [56, 13, 71], symbol tuning [100], black-box tuning [86], reinforcement learning from human feedback [78], chain-of-thought reasoning [102, 52, 111, 117], and more.
Recent relevant research includes works on tensor transformers [83, 4, 69, 58, 116], acceleration techniques [103, 101, 89, 61, 81, 92, 39, 90, 84, 99, 50, 40, 42, 60, 47, 14, 51, 62, 59, 46, 67, 65, 48, 64, 3, 18], and other related studies [25, 87, 36, 68, 91, 15, 105, 70, 27, 33, 17, 57, 44, 88, 119, 112, 63].
Appendix B Tools From Previous Work
Here, we present the lemma of Iteratively visiting to find minimum value.
Lemma B.1 (Get minimum value, Implicitly in [24]).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.10.
Appendix C Missing Proof in Key Operation Implementation
@@
In Section C.1, the operation of selection is discussed. In Section C.2, the operation of increment is discussed. In Section C.4, the operation of reading scalar from column is discussed. In Section C.3, the operation of comparison is discussed. In Section C.5, the operation of writing scalar to column is discussed. In Section C.6, the operation of termination is discussed. In Section C.7, the operation of reading from incident matrix is discussed. In Section C.8, the operation of AND is discussed. In Section C.9, the operation of repeat AND is discussed. In Section C.10, the operation of repeat addition is discussed.
C.1 Selection
Lemma C.1 (Selection, formal version of Lemma C.1).
@@ C.1 Selection
C.2 Increment
Lemma C.2 (Increment, formal version of Lemma C.2).
@@ C.2 Increment
  • Let $C_{1},C_{2}$ be the index for source position embedding, $D_{1},D_{2}$ be the index for target position embedding.
@@
C.3 Comparison
Lemma C.3 (Comparison, formal version of Lemma C.3).
@@ C.3 Comparison
  • Let $C,~{}D$ be the index for the column for comparing.
  • Let $E$ be the index for the target column to write the comparison result.
Proof.
C.4 Read Scalar From Column
Lemma C.4 (Read scalar from column, formal version of Lemma C.4).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.10.
  • Let $\Omega$ represent the maximum absolute value within a clause.
  • Let $C_{1},C_{2}$ be the position embedding for source row, $C$ be the row index for the source scalar, and $D$ be the cloumn index for source scalar.
@@ C.4 Read Scalar From Column
C.5 Write Scalar to Column
Lemma C.5 (Write scalar to column, formal version of Lemma C.5).
@@ C.5 Write Scalar to Column
C.6 Termination
Lemma C.6 (Termination, formal version of Lemma C.6).
@@ C.6 Termination
C.7 Read from Incident Matrix
Lemma C.7 (Read from incident matrix, formal version of Lemma C.7).
@@ C.8 AND
C.9 Repeat AND
Lemma C.9 (Repeat AND, formal version of Lemma C.9).
@@ C.9 Repeat AND
C.10 Repeat Addition
Lemma C.10 (Repeat addition, formal version of Lemma C.10).
@@ C.10 Repeat Addition
Algorithm 3 Iteration of minimum value
3:$\mathrm{visit}_{\min}\leftarrow{\bf 0}_{n_{v}}\|{\bf 1}_{K-1-n_{v}}$
5:
6:procedure GetMinimumValue($x\in\mathbb{R}^{d}$)
7:     if $\mathrm{termination}_{\min}=1$ then
8:         $\cdots$ $\triangleright$ Re-initialization of minimum value, Lemma C.1
9:         $\mathrm{termination}_{\min}\leftarrow 0$
10:     end if
11:     $\mathrm{idx}_{\mathrm{cur}}\leftarrow\mathrm{idx}_{\mathrm{cur}}+1$triangleright Increment, Lemma C.2
12:     $\mathrm{val}_{\mathrm{cur}}\leftarrow x[\mathrm{idx}_{\mathrm{cur}}]$triangleright Read scalar from column, Lemma C.4
13:     if $\mathrm{val}_{\mathrm{cur}}<\mathrm{val}_{\mathrm{best}}$ $\triangleright$ Compare, Lemma C.3
14:         $\mathrm{val}_{\mathrm{best}},~{}\mathrm{idx}_{\mathrm{best}}\leftarrow\mathrm{val}_{\mathrm{cur}},~{}\mathrm{idx}_{\mathrm{cur}}$triangleright Update variables, Lemma C.1
15:     end if
16:     $\mathrm{visit}_{\min}[\mathrm{idx}_{\mathrm{cur}}]\leftarrow 1$triangleright Write scalar to column, Lemma C.5
17:     $\mathrm{termination}_{\min}\leftarrow\neg(0~{}\mathrm{in}~{}\mathrm{visit}_{\min})$triangleright Trigger termination, Lemma C.6
18:end procedure
Algorithm 4 Dijkstra’s algorithm for shortest path
5:
6:    while $\mathrm{termination}$ is $\mathbf{false}$ do
7:         for $i=1\to K-1$ do
8:             if $\mathrm{visit}[i]$ is $\mathbf{true}$ then
9:                 $\mathrm{dists}_{\mathrm{masked}}[i]\leftarrow\Omega$triangleright Selection, Lemma C.1
10:             else
11:                 $\mathrm{dists}_{\mathrm{masked}}[i]\leftarrow\mathrm{dists}[i]$
12:             end if
13:         end for
14:
15:         GetMinimumValue $(\mathrm{dists}_{\mathrm{masked}})$triangleright Get minimum value, Lemma B.1
16:
17:         if $\mathrm{termination}_{\min}$ is $\mathbf{true}$ then
18:             $\mathrm{node}\leftarrow\mathrm{idx}_{\mathrm{best}}$
19:             $\operatorname{dist}\leftarrow\mathrm{val}_{\mathrm{best}}$triangleright Selection, Lemma C.1
20:         end if
21:
22:         VisitHyperdege $(A)$triangleright Degradation, Theorem 4.1
23:
24:         if $\mathrm{termination}_{\mathrm{hyperedge}}=1$ then
25:             $\mathrm{candidates}\leftarrow\mathrm{candidates}_{\mathrm{hyperedge}}$triangleright Selection, Lemma C.1
26:         end if
27:         for $i=1\to K-1$ do
28:             $\mathrm{candidates}[i]\leftarrow\mathrm{candidates}[i]+\operatorname{dist}$triangleright Addition, Lemma C.10
29:         end for
30:         for $i=1\to K-1$ do
31:             $\mathrm{changes}[i]\leftarrow\mathrm{candidates}[i]<\mathrm{dists}[i]$triangleright Compare, Lemma C.3
32:         end for
33:         for $i=1\to K-1$ do
34:             if $\mathrm{termination}_{\min}=0$ and $\mathrm{iszero}[i]=1$ then
35:                 $\mathrm{changes}[i]\leftarrow 0$triangleright repeat AND, Lemma C.9
36:             end if
37:         end for
38:         for $i=1\to K-1$ do
39:             if $\mathrm{changes}[i]=1$ then
40:                 $\mathrm{prev}[i],~{}\mathrm{dists}[i]\leftarrow\mathrm{node},\mathrm{candidates}[i]$triangleright Selection. Lemma C.1
41:             end if
42:         end for
43:         $\mathrm{visit}[\mathrm{node}]\leftarrow\mathrm{visit}[\mathrm{node}]+\mathrm{termination}_{\min}$triangleright Write scalar to column, Lemma C.5
44:         $\mathrm{termination}\leftarrow\neg(0~{}\mathrm{in}~{}\mathrm{visit})$triangleright Trigger termination, Lemma C.6
45:    end while
46:    return $\mathrm{prev},~{}\mathrm{dists}$
47:end procedure
Algorithm 5 Helly, Algorithm 3 in [10]
1:$\mathrm{idx}_{x}\leftarrow 0$
2:$\mathrm{idx}_{y}\leftarrow 1$
3:$\mathrm{idx}_{v}\leftarrow 0$
4:$\mathrm{termination}\leftarrow 0$
5:$\mathrm{helly}\leftarrow 1$triangleright Initialization of Helly
6:
7:while $\mathrm{termination}=0$ do
8:     $\mathrm{hyperedge}_{x}\leftarrow A[:,\mathrm{idx}_{x}]$
9:     $\mathrm{hyperedge}_{y}\leftarrow A[:,\mathrm{idx}_{y}]$
10:     $\mathrm{hyperedge}_{v}\leftarrow A[:,\mathrm{idx}_{v}]$triangleright Read column from incident matrix, Lemma C.7
11:     $\mathrm{hyperedge}_{x}\leftarrow\mathrm{hyperedge}_{x}>0$
12:     $\mathrm{hyperedge}_{y}\leftarrow\mathrm{hyperedge}_{y}>0$
13:     $\mathrm{hyperedge}_{v}\leftarrow\mathrm{hyperedge}_{v}>0$triangleright Compare, Lemma C.3
14:     $\mathrm{intersection}\leftarrow\mathrm{hyperedge}_{x}\wedge\mathrm{hyperedge}_{y}\wedge\mathrm{hyperedge}_{v}$triangleright AND, Lemma C.8
15:     $\mathrm{helly}_{v}\leftarrow\neg$(1 in $\mathrm{intersection}$)
16:     if $\mathrm{helly}_{v}=1$ then
17:         $\mathrm{helly}\leftarrow 0$
18:         $\mathrm{termination}\leftarrow 1$triangleright Selection, Lemma C.1
19:     end if
20:     $\mathrm{idx}_{v}\leftarrow\mathrm{idx}_{v}+1$triangleright Increment, Lemma C.2
21:     if $\mathrm{idx}_{v}>n_{v}$ then
22:         $\mathrm{idx}_{v}\leftarrow 1$triangleright Selection, Lemma C.1
23:         $\mathrm{idx}_{y}\leftarrow\mathrm{idx}_{y}+1$triangleright Selection and Increment, Lemma C.1 and Lemma C.2
24:     end if
25:     if $\mathrm{idx}_{y}>n_{v}$ then
26:         $\mathrm{idx}_{x}\leftarrow\mathrm{idx}_{x}+1$
27:         $\mathrm{idx}_{y}\leftarrow\mathrm{idx}_{x}+1$triangleright Selection and Increment, Lemma C.1 and Lemma C.2
28:     end if
29:     if $\mathrm{idx}_{x}=n_{v}$ then
30:         $\mathrm{termination}\leftarrow 1$triangleright Selection, Lemma C.1
31:     end if
32:end while
33:return $\mathrm{helly}$
Appendix D Missing Proof in Simulation
Theorem D.1 (Dijkstra’s Algorithm on hypergraph, formal version of).
Proof.
Let Dijkstra’s algorithm be considered as described in Algorithm 4. Following from Lemma 6.1 and Lemma B.1, the operation of iteratively visiting vertices and hyperedges requires 18 layers, as established in these lemmas. For the remaining part of the algorithm, the operations include 4 selection operations, 1 add operation, 1 compare operation, 1 repeat AND operation, 1 write scalar to column operation, and 1 trigger termination operation. According to Lemma C.1, Lemma C.10, Lemma C.9, Lemma C.3, Lemma C.5, and Lemma C.6, these operations together require 9 layers. Therefore, the total number of layers required to simulate Dijkstra’s algorithm is:
$18+9=27$
layers of the transformer. ∎
Lemma D.2 (Visiting hyperedges iteratively, formal version of Lemma 6.1).